Prove that electric field and electric potential are connected by the formula E=V/r

1
Expert's answer
2021-11-22T10:13:02-0500

The potential of an electric field of a point charge is given by

"V=k\frac{q}{r}"

The strength of an electric field of a point charge is given by

"E=k\frac{q}{r^2}"

Hence

"E=\frac{V}{r}"